

********************************************************************
UARS WEEKLY STATUS REPORT - 13 Aug 2004
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to perform nominally, taking solar data on all
available orbital tracking periods. This week we received about 59% of
each day's data amounting to nearly 547 minutes of solar observing time
per day.
***********
SUSIM
***********
SUSIM successfully gathered its daily mid (1.1 nm) resolution, MgII,
occultation, solar pointing offset, and continuous monitoring scans
for periods of UARS real-time telemetry or onboard tape recording.
However, because of the unexpected absence of TDRS telemetry on
August 10, the half of the daily mid-resolution scans were lost on
that day.
***********
PEM
***********
HEPSA data files are version 2. Currently, HEPSA data between 1991274
and 2003365 has been reprocessed and archived at the GSFC DAAC. Currently,
HEPSA data between 2004001 and 2004039 has been reprocessed waiting for
archival at the GSFC DAAC.
MEPS data files are version 1. Processed MEPS data files at the
DAAC archive are between 1991274 and 2004365. Continued processing of MEPS
data has restarted.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html. The additional electron data for
the October and November storms of 2003 has been moved and is now located at
http://climatology.space.swri.edu/uars_fromRudy.html. For those users of
SDDAS, update your UARS Meta data and these data times will appear
with full functionality. Proton data from these storms are now located at
http://climatology.space.swri.edu/UARS_images_from_Rudy_(Apr._2004).html
Requests have been received for storm spectral data. Overview plots were
produced and these plots may be found on the web at
http://climatology.space.swri.edu/UARS_images_postedJune2004.html
for the storms of October 29 of 1991, November 8-9 of 1991, April 6-7 of 2000,
March 31 - April 1 of 2001, April 11-12 of 2001, November 6-7 of 2001, and
November 11-12 of 2001.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 31 July 2004 (Orbit 70518, GMT Day 213) through
06 August 2004 (Orbit 70623, GMT Day 219).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 218/1446 70602 POWER OFF
The beta angle ranged between the angles of 37.9 degrees to 36.7
degrees for this report period. The beta angle is now decreasing
towards a minimum angle of 0 degrees on 18 August 2004 (DOY 231).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 0.95 to 1.28 degrees C. Additional
battery performance data for each battery on 30 July and 06 August
2004 are provided in Appendix D.
The clock error ranged from 11.9 to -11.4 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 03
HRDI = 01
PEM = 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 03
Three (3) Anomaly Reports remained OPEN at the end of this report
period (see Appendix E).
GROUND SYSTEM OPERATION
Three (3) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
This week, between 02 and 05 August, the FOT observed an increase in
the battery 2 differential voltage. The differential voltage went
from its nominal value of -40 to-80 mV to -230 mV. This is the third
instance of this since May of this year. In order to alleviate this
problem the FOT powered off PEM Axis nine hours early on DOY 218 and
eliminated the HRDI to fulltime period from DOY 219-224. In the
future the FOT will go to VT level 6 as the first course of action and
then reduce load if necessary.
The FOT completed the testing of the new version of the Eclipse ground
system software. Only one out of the fourteen DRs delivered did not
function properly. The build will be released to operations despite
this discrepancy. Due to scheduling issues with key personnel during
the next few weeks, the build will not be put onto the servers until
07 September 2004.
The new CMS code to allow for SMA events continued to undergo testing
this week. Problems receiving products from FDF, transferring the
schedule from UPS to CMS, and system drive issues hindered progress.
While the FDF product and system drive issues have been resolved, the
transfer issue is still under investigation.
MEETINGS
The FOT attended the UARS Monthly Status review on 03 August 2004, to
discuss the significant events from the month of July.
The FOT called a meeting on 06 August 2004 to discuss the increase in
the differential voltage and how to manage future occurrences. The
results of this meeting were that everyone agreed that this is a sign
of ageing of the battery, and in the future VT-6 will be utilized as
the first recourse to alleviate the problem.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 18 August 2004 (DOY 231).
This maneuver will be a Reverse to Forward Flight maneuver (Southern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 06 August 2004
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
70524 213/0916 04-225 OBC/CLK CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
70526 213/1344 04-226 OBC/CLK CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
70546 214/2032 04-227 OBC/CLK CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
70602 218/1446 04-228 PEM AXIS POWER OFF
70606 218/2056 04-229 HRDI HRDI MICROLOADS & NOOPS
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 06 August 2004
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 06 August 2004
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
70524 213/0936 CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
70526 213/1328 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
70546 214/2039 CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 06 August 2004
30 July 2004 - GMT Day 212 Beta = 36.8 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.05/-4.31 +00.0/+00.0 07.0 0.00 0.0
2 -00.0/-184.8 +5.15/+4.51 +26.4/-14.4 25.6 1.02 20.0
3 +33.6/-072.8 +4.51/+3.23 +25.6/-14.4 25.6 1.02 19.8
06 August 2004 - GMT Day 219 Beta = 36.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.05/-4.31 +00.0/+00.0 07.0 0.00 0.0
2 -00.0/-212.8 +5.15/+3.87 +26.8/-14.4 25.6 1.04 20.0
3 +33.6/-033.6 +3.87/+2.92 +26.4/-14.4 25.4 1.03 19.5
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
-None this report period.
APPENDIX E
UARS ANOMALY REPORTS
Week ending 06 August 2004
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-050 FOT 70538 214/0729 30832 Generic Late Acquisition
04-051 FOT 70564 216/0211 36541 Data Dropout
04-052 FOT 70576 216/2103 36556 Loss of Downlink
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-039 FOT See List July 04 30832 Generic Late Acquisition
04-049 FOT 70501 211/2042 36507 Data Dropout - weather W/S
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-050 FOT 70538 214/0729 30832 Generic Late Acquisition
04-051 FOT 70564 216/0211 36541 Data Dropout
04-052 FOT 70576 216/2103 36556 Loss of Downlink